# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 5.2. Arithmetic compactifications of splitting models
We very briefly review some basic properties of the toroidal and minimal compactifications of $\mathsf{M}_{K}^{\mathrm{spl}}$ constructed by Lan in [35], which will be used in the following.
Recall that we have a morphism $\mathsf{M}_{K}^{\mathrm{spl}}\rightarrow\mathsf{M}_{K}\otimes\operatorname{\mathcal{O}}_{F}$. By [33, 34] we choose a toroidal compactification $\mathsf{M}^{\mathrm{tor}}_{K,\Sigma}$ of $\mathsf{M}_{K}$ over $\operatorname{\mathcal{O}}_{E}$ associated to a compatible collection $\Sigma$ of cone decompositions. Over $\mathsf{M}_{K}$, we have the polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{\mathsf{M}_{K}}$-modules (¯,¯$(\underline{\mathcal{H}},\underline{\mathscr{F}},\underline{\iota})$ associated to the universal abelian scheme with additional structure as above. By [35] Proposition 3.1.2, the triple (¯,¯$(\underline{\mathcal{H}},\underline{\mathscr{F}},\underline{\iota})$ uniquely extends to an $\operatorname{\mathscr{L}}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{\mathsf{M}^{\mathrm{tor}}_{K,\Sigma}}$-modules (¯$(\underline{\mathcal{H}}^{\mathrm{ext}},\underline{\operatorname{\mathscr{F}}}^{\mathrm{ext}},\underline{\iota}^{\mathrm{ext}})$. Then we define
@@
as the scheme of splitting structures of (¯$(\underline{\mathcal{H}}^{\mathrm{ext}},\underline{\operatorname{\mathscr{F}}}^{\mathrm{ext}},\underline{\iota}^{\mathrm{ext}})$ over $\mathsf{M}^{\mathrm{tor}}_{K,\Sigma}\otimes\operatorname{\mathcal{O}}_{F}$. By [35] Theorem 3.4.1, for suitable choice of $\Sigma$, the scheme $\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$ is normal, projective and flat over $\operatorname{\mathcal{O}}_{F}$, which admits a similar description as the usual compactifications $\mathsf{M}^{\mathrm{tor}}_{K,\Sigma}$.
@@
By [35] Propositions 4.1.22, 4.2.31 and 4.2.34, we have the minimal compactification $\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$ together with a canonical morphism $\oint:\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}\longrightarrow\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$, which fits into a commutative diagram of schemes over $\operatorname{\mathcal{O}}_{F}$:
@@
By loc. cit. Theorem 4.3.1, the minimal compactification $\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$ admits a similar description as $\mathsf{M}_{K}^{\mathrm{min}}$.
Let $Z^{\mathrm{spl}}\subset\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$ be a boundary stratum. There is a corresponding boundary stratum $Z\subset\mathsf{M}_{K}^{\mathrm{min}}$, which is an analogue of $\mathsf{M}_{K}$ for some boundary PEL datum $(\operatorname{\mathcal{O}}_{B},\ast,\Lambda^{Z},\langle\cdot,\cdot\rangle^{Z},h^{Z})$ so that we have the tautological abelian scheme $(B,\lambda,\iota)$ over $Z$ with associated (¯,¯$({}^{\sharp}\underline{\mathcal{H}},{}^{\sharp}\underline{\operatorname{\mathscr{F}}},{}^{\sharp}\underline{\iota})$. By construction, we have
@@
  • an admissible cone decomposition $\Sigma_{Z}$ of some $P=P_{Z}$, as well as a subset $\Sigma_{Z}^{+}$ of $\Sigma_{Z}$ which forms a cone decomposition of the interior $P^{+}$ of $P$
  • an arithmetic group $\Gamma=\Gamma_{Z}$ acting on $P$ and hence aslo on $\Sigma_{Z}$; the open cone $P^{+}$ and the corresponding $\Sigma_{Z}^{+}$ are stable under the action of $\Gamma$. As in [36] we may and we shall assume that for each $\sigma\in\Sigma_{Z}^{+}$, the stabilizer $\Gamma_{\sigma}$ is trivial,
  • a finite free abelian group $S$; let $E$ be a split torus over $\mathbb{Z}$ with character group $S$.
@@
  • a morphism of schemes $\Xi^{\mathrm{spl}}\rightarrow C^{\mathrm{spl}}$ which is an $E$-torsor; for each $\sigma\in\Sigma_{Z}$, we have an affine toroidal embedding $\Xi^{\mathrm{spl}}\hookrightarrow\Xi^{\mathrm{spl}}(\sigma)$ over $C^{\mathrm{spl}}$ with a closed subscheme $\Xi_{\sigma}^{\mathrm{spl}}$.
  • for each representative $\sigma\in\Sigma_{Z}^{+}$ of an orbit $[\sigma]\in\Sigma_{Z}^{+}/\Gamma$, let $Z_{[\sigma]}^{\mathrm{spl}}\subset\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$ be the corresponding toroidal boundary stratum, and $\mathfrak{X}^{\mathrm{spl}}_{\sigma}:=(\Xi^{\mathrm{spl}}(\sigma))^{\wedge}_{\Xi^{\mathrm{spl}}_{\sigma}}$ the formal completion, then there is a canonical isomorphism of formal schemes
@@
    $\mathfrak{X}^{\mathrm{spl}}:=\mathfrak{X}_{Z}^{\mathrm{spl}}=(\Xi_{Z}^{\mathrm{spl}})^{\wedge}_{\cup_{\tau\in\Sigma_{Z}^{+}}\Xi_{\tau}^{\mathrm{spl}}}$
@@
    $\mathfrak{X}^{\mathrm{spl}}/\Gamma\simeq(\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}})^{\wedge}_{\cup_{\tau\in\Sigma_{Z}^{+}/\Gamma}Z_{[\tau]}^{\mathrm{spl}}}.$
@@
  • there are similar and parallel objects $C,\Xi$ etc for the boundary stratum $Z$, such that $C^{\mathrm{spl}}=\textrm{Spl}^{+}_{({}^{\sharp}\underline{\mathcal{H}},{}^{\sharp}\underline{\operatorname{\mathscr{F}}},{}^{\sharp}\underline{\iota})/C\otimes\operatorname{\mathcal{O}}_{F}},\quad\Xi^{\mathrm{spl}}=\textrm{Spl}^{+}_{({}^{\sharp}\underline{\mathcal{H}},{}^{\sharp}\underline{\operatorname{\mathscr{F}}},{}^{\sharp}\underline{\iota})/\Xi\otimes\operatorname{\mathcal{O}}_{F}}$ etc, see [35] Lemma 3.2.4.
@@ 5.3. Canonical extensions of automorphic vector bundles
Since $G$ is split over $F$, it defines a reductive group scheme over $\operatorname{\mathcal{O}}_{F}$ which is what we denoted by $\mathcal{G}^{\mathrm{spl}}$. Moreover, the parabolic subgroup $P_{\mu}$ extends to a paraholic group scheme of $\mathcal{G}^{\mathrm{spl}}$ over $\operatorname{\mathcal{O}}_{F}$. Thus the flag variety $\mathscr{F}\ell(G,\mu)_{F}$ extends canonically to a smooth scheme $\mathcal{G}^{\mathrm{spl}}/P_{\mu}$ (the integral flag variety) over $\operatorname{\mathcal{O}}_{F}$ which is nothing else but $\prod_{i\in I,1\leq j\leq f_{i},1\leq l\leq e_{i}}\mathbb{M}_{i,j}^{l}$. We still denote by $\mathscr{F}\ell(G,\mu)$ the integral flag variety over $\operatorname{\mathcal{O}}_{F}$.
@@
where $\pi$ is a $\operatorname{\mathcal{G}}^{\mathrm{spl}}$-torsor and $q$ is $\operatorname{\mathcal{G}}^{\mathrm{spl}}$-equivariant. For any representation $(V,\eta)\in\mathrm{Rep}_{\operatorname{\mathcal{O}}_{F}}P_{\mu}$, we get the associated $G$-equivariant vector bundle $\underline{V}$ on $\mathscr{F}\ell(G,\mu)$. Via the above diagram, we get a vector bundle $\mathcal{V}^{\mathrm{can}}=\mathcal{V}_{\eta}^{\mathrm{can}}$ on $\mathsf{M}^{\mathrm{spl},\mathrm{tor}}_{K,\Sigma}$, which we call the canonical extension of the vector bundle $\mathcal{V}=\mathcal{V}_{\eta}$ on $\mathsf{M}_{K}^{\mathrm{spl}}$. Let $D^{\prime}$ be the effective Cartier divisor of [35] Corollary 4.4.4 with $D^{\prime}_{red}=\mathsf{M}^{\mathrm{spl},\mathrm{tor}}_{K,\Sigma}\setminus\mathsf{M}_{K}^{\mathrm{spl}}$ and set $\mathcal{V}^{\mathrm{sub}}=\mathcal{V}^{\mathrm{can}}(-D^{\prime})$, which we call the sub canonical extension of $\mathcal{V}=\mathcal{V}_{\eta}$.
@@
    ^$\widehat{\bigoplus}_{\ell\in\sigma^{\vee}}\Big{(}(\Psi^{\mathrm{spl}}_{Z}(\ell))^{\wedge}_{\overline{x}}\otimes\mathcal{V}_{0,\overline{x}}\Big{)},$
@@
One needs to check in the smooth reduction case the sheaf $\mathcal{V}^{\mathrm{can}}$ satisfies the above two conditions. This can be achieved by modifying the arguments in the proof of [32] Proposition 5.6, where we take account of the splitting structures everywhere. ∎
@@ 5.4. Extensions of $\mathcal{G}_{0}^{\mathrm{spl}}$-zips and Hasse invariants
Before discussing extensions of zips to the boundary, we need some group theoretical preparation. For each $Z^{\mathrm{spl}}\subset\mathsf{M}^{\mathrm{spl},\mathrm{min}}_{K,0}$, recall that we have the boundary PEL-type $\operatorname{\mathcal{O}}_{B}$-lattice $(\Lambda^{Z},\langle\cdot,\cdot\rangle^{Z},h^{Z})$. Over $\mathbb{Q}$, we have the associated parabolic subgroups
$Q=Q_{Z}\subset P=P_{Z}\subset G$
with the same unipotent radical $U$ such that $M_{h}:=Q/U$ is the reductive group for the rational PEL datum corresponding to $Z$, and $M:=P/U=M_{h}\times M_{l}$ where $M_{l}\simeq P/Q$ is a reductive group factor of the Levi $M$, see for example [36] Definition 3.8. Consider the reductive group $\mathcal{G}_{0}^{\mathrm{spl}}$ over $\operatorname{\mathbb{F}}_{p}$. We want to adapt the construction of [2] subsection 3.3 to our setting. Consider the induced symplectic filtration
@@
For (1), we just repeat the construction in subsection 3.2 starting from (¯$(\underline{\mathcal{H}}^{\mathrm{ext}},\underline{\operatorname{\mathscr{F}}}^{\mathrm{ext}},\underline{\iota}^{\mathrm{ext}})$ together with its universal splitting structure.
@@
is smooth. If $x\in\mathsf{M}^{\mathrm{spl}}_{K,0}$, this has been done in Proposition 3.10. So we may assume $x\in Z_{[\sigma]}^{\mathrm{spl}}$ for a boundary stratum. Then there is corresponding point $x^{\prime}\in\Xi^{\mathrm{spl}}(\sigma)$ such that $C=\widehat{\operatorname{\mathcal{O}}}_{\Xi^{\mathrm{spl}}(\sigma),x^{\prime}}$. Let $y\in Z^{\mathrm{spl}}$ be the image of $x$ and $D=\widehat{\operatorname{\mathcal{O}}}_{Z^{\mathrm{spl}},y}$. We get an induced morphism
@@
To show the smoothness of $\zeta_{C}:\operatorname{Spec}\,C\rightarrow\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{\kappa}$, we proceed in four steps. Step 1. We show $\zeta_{C}=\tau_{Z}\circ\zeta_{D}\circ\oint$. In other words, the following diagram commutes
@@
We need to analysis the structure of the $\mathcal{G}_{0}^{\mathrm{spl}}$-zip on $C$ in terms of the associated $\mathcal{G}_{0,Z}^{\mathrm{spl}}$-zip. Consider the restriction of (¯$(\underline{\mathcal{H}}^{\mathrm{ext}},\underline{\operatorname{\mathscr{F}}}^{\mathrm{ext}},\underline{\iota}^{\mathrm{ext}})$ on $Z_{[\sigma]}^{\mathrm{spl}}$, which we denote by (¯,¯$(\underline{\mathcal{H}}^{\natural},\underline{\operatorname{\mathscr{F}}}^{\natural},\underline{\iota}^{\natural})$. By [35] proof of Proposition 3.3.21, there are isomorphisms
@@
Moreover, the filtration $\underline{\operatorname{\mathscr{F}}_{\bullet}}^{\natural}$ on $\operatorname{\mathscr{F}}^{\natural}$ induces filtrations on ${}^{\sharp}\operatorname{\mathscr{F}}$ and ${}^{\flat}\operatorname{\mathscr{F}}$ as [35] Lemma 3.3.11, and by loc. cit. Corollary 3.3.16 the filtration on ${}^{\flat}\operatorname{\mathscr{F}}$ is independent of the filtration on $\operatorname{\mathscr{F}}^{\natural}$. By the construction in section 3, the $F$-zip $\mathcal{M}^{\natural}$ (constructed from (¯,¯$(\underline{\mathcal{H}}^{\natural},\underline{\operatorname{\mathscr{F}}}^{\natural},\underline{\iota}^{\natural})$) is a direct sum of the $F$-zip $\mathcal{M}_{Z}$ constructed from (¯,¯$({}^{\sharp}\underline{\mathcal{H}},{}^{\sharp}\underline{\operatorname{\mathscr{F}}},{}^{\sharp}\underline{\iota})$ and the $F$-zip constructed from ${}^{\flat}\operatorname{\mathscr{F}}$. Translating further into the language of $\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}$-zips, we have the factorization of $\zeta_{C}$. Step 2. $\oint:\operatorname{Spec}\,C\rightarrow\operatorname{Spec}\,D$ is smooth. For this, we apply the local diagrams for $\mathsf{M}^{\mathrm{spl},\mathrm{tor}}_{K,\Sigma}$ and $Z^{\mathrm{spl}}$ to realize $C$ and $D$ as local rings of the corresponding local models, which are also completions along identities of some unipotent subgroups, see the proof of Proposition 3.10. Then one can prove the smoothness of $\oint$ by checking that on tangent spaces it induces a projection. Step 3. $\zeta_{D}:\operatorname{Spec}\,D\rightarrow\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}$ is smooth. As $Z^{\mathrm{spl}}$ is smooth, the smoothness of $\zeta_{D}$ follows from the boundary version of Proposition 3.10. Step 4. $\tau_{Z}:\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}\rightarrow\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}$ is smooth. This follows from a similar argument as the proof of [2] Lemma 3.4. ∎
@@
The Hasse invariants $h_{w}\in H^{0}(\overline{X_{w}},\omega_{\operatorname{Hdg}}^{N_{w}})$ of Corollary 4.7 extends to a $G(\operatorname{\mathbb{A}}_{f}^{p})$-equivariant section $h_{w}^{\mathrm{tor}}\in H^{0}(\overline{X^{\mathrm{tor}}_{w}},\omega_{\operatorname{Hdg}}^{N_{w}})$ with non vanishing locus precisely $X_{w}^{\mathrm{tor}}$.
@@
Similar to [2] Corollary 3.7, this follows from the fact that for each boundary $Z$, $\tau_{Z}:\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}\rightarrow\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}$ is smooth, thus open, therefore does not contain the closed point of $\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}$. ∎

### v3

@@ 5.2. Arithmetic compactifications of splitting models
We briefly review some basic properties of the toroidal and minimal compactifications of $\mathsf{M}_{K}^{\mathrm{spl}}$ constructed by Lan in [37], which will be used in the following.
[…]
@@
as the scheme of splitting structures of $(\underline{\mathcal{H}}^{\mathrm{ext}},\underline{\operatorname{\mathscr{F}}}^{\mathrm{ext}},\underline{\iota}^{\mathrm{ext}})$ over $\mathsf{M}^{\mathrm{tor}}_{K,\Sigma}\otimes\operatorname{\mathcal{O}}_{F}$. By [37] Theorem 3.4.1, for suitable choice of $\Sigma$, the scheme $\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$ is normal, projective and flat over $\operatorname{\mathcal{O}}_{F}$, and admits a similar description as the usual compactifications $\mathsf{M}^{\mathrm{tor}}_{K,\Sigma}$.
@@
By [37] Propositions 4.1.22, 4.2.31 and 4.2.34, we have the minimal compactification $\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$ together with a canonical morphism $\oint:\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}\rightarrow\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$, which fits into a commutative diagram of schemes over $\operatorname{\mathcal{O}}_{F}$:
@@
By loc. cit. Theorem 4.3.1, the minimal compactification $\mathsf{M}_{K}^{\mathrm{spl},\mathrm{min}}$ admits a similar description to $\mathsf{M}_{K}^{\mathrm{min}}$.
[…]
@@
  • an admissible cone decomposition $\Sigma_{Z}$ of some cone $\mathbf{P}=\mathbf{P}_{Z}$, as well as a subset $\Sigma_{Z}^{+}$ of $\Sigma_{Z}$ which forms a cone decomposition of the interior $\mathbf{P}^{+}$ of $\mathbf{P}$.
  • an arithmetic group $\Gamma=\Gamma_{Z}$ acting on $\mathbf{P}$ and hence aslo on $\Sigma_{Z}$; the open cone $\mathbf{P}^{+}$ and the corresponding $\Sigma_{Z}^{+}$ are stable under the action of $\Gamma$. As in [38] we may and we shall assume that for each $\sigma\in\Sigma_{Z}^{+}$, the stabilizer $\Gamma_{\sigma}$ is trivial.
  • a finite free abelian group $S$; let $\mathbf{E}=\mathbf{E}_{Z}$ be a split torus over $\mathbb{Z}$ with character group $S$.
@@
  • a morphism of schemes $\Xi^{\mathrm{spl}}\rightarrow C^{\mathrm{spl}}$ which is an $\mathbf{E}$-torsor; for each $\sigma\in\Sigma_{Z}$, we have an affine toroidal embedding $\Xi^{\mathrm{spl}}\hookrightarrow\Xi^{\mathrm{spl}}(\sigma)$ over $C^{\mathrm{spl}}$ with a closed subscheme $\Xi_{\sigma}^{\mathrm{spl}}$.
  • for each representative $\sigma\in\Sigma_{Z}^{+}$ of an orbit $\sigma]\in\Sigma_{Z}^{+}/\Gamma$, let $Z_{[\sigma]}^{\mathrm{spl}}\subset\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$ be the corresponding toroidal boundary stratum, and $\mathfrak{X}^{\mathrm{spl}}_{\sigma}:=(\Xi^{\mathrm{spl}}(\sigma))^{\wedge}_{\Xi^{\mathrm{spl}}_{\sigma}}$ the formal completion, then there is a canonical isomorphism of formal schemes
@@
    $\mathfrak{X}^{\mathrm{spl}}:=\mathfrak{X}_{Z}^{\mathrm{spl}}=(\Xi_{Z}^{\mathrm{spl}})^{\wedge}_{\quad\cup_{\tau\in\Sigma_{Z}^{+}}\Xi_{\tau}^{\mathrm{spl}}}$
@@
    $\mathfrak{X}^{\mathrm{spl}}/\Gamma\simeq(\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}})^{\wedge}_{\quad\cup_{\tau\in\Sigma_{Z}^{+}/\Gamma}Z_{[\tau]}^{\mathrm{spl}}}.$
@@
  • there are similar and parallel objects $C,\Xi$, etc. for the boundary stratum $Z$, such that $C^{\mathrm{spl}}=\textrm{Spl}^{+}_{({}^{\sharp}\underline{\mathcal{H}},{}^{\sharp}\underline{\operatorname{\mathscr{F}}},{}^{\sharp}\underline{\iota})/C\otimes\operatorname{\mathcal{O}}_{F}},\quad\Xi^{\mathrm{spl}}=\textrm{Spl}^{+}_{({}^{\sharp}\underline{\mathcal{H}},{}^{\sharp}\underline{\operatorname{\mathscr{F}}},{}^{\sharp}\underline{\iota})/\Xi\otimes\operatorname{\mathcal{O}}_{F}}$, etc., see [37] Lemma 3.2.4.
@@ 5.3. Canonical extensions of automorphic vector bundles
Since $G$ is split over $F$, it defines a reductive group scheme over $\operatorname{\mathcal{O}}_{F}$, denoted by $\mathcal{G}^{\mathrm{spl}}$ (or $G$) as before. Moreover, recall that as in the paragraph above Corollary 2.4 the parabolic subgroup $P_{\mu}$ extends to a paraholic group scheme of $\mathcal{G}^{\mathrm{spl}}$ over $\operatorname{\mathcal{O}}_{F}$. Thus the flag variety $\mathscr{F}\ell(G,\mu)_{F}$ extends canonically to a smooth scheme $\mathcal{G}^{\mathrm{spl}}/P_{\mu}$ (the integral flag variety) over $\operatorname{\mathcal{O}}_{F}$ which is nothing else but $\prod_{i\in I,1\leq j\leq f_{i},1\leq l\leq e_{i}}\mathbb{M}_{i,j}^{l}$. We still denote by $\mathscr{F}\ell(G,\mu)$ the integral flag variety over $\operatorname{\mathcal{O}}_{F}$.
@@
where $\pi$ is a $\operatorname{\mathcal{G}}^{\mathrm{spl}}$-torsor and $q$ is $\operatorname{\mathcal{G}}^{\mathrm{spl}}$-equivariant. For any representation $(V,\eta)\in\mathrm{Rep}_{\operatorname{\mathcal{O}}_{F}}P_{\mu}$, we get the associated $G$-equivariant vector bundle $\underline{V}$ on $\mathscr{F}\ell(G,\mu)$. Via the above diagram, we get a vector bundle $\mathcal{V}^{\mathrm{can}}=\mathcal{V}_{\eta}^{\mathrm{can}}$ on $\mathsf{M}^{\mathrm{spl},\mathrm{tor}}_{K,\Sigma}$, which we call the canonical extension of the vector bundle $\mathcal{V}=\mathcal{V}_{\eta}$ on $\mathsf{M}_{K}^{\mathrm{spl}}$. Let $D^{\prime}$ be the effective Cartier divisor of [37] Corollary 4.4.4 with $D^{\prime}_{red}=\mathsf{M}^{\mathrm{spl},\mathrm{tor}}_{K,\Sigma}\setminus\mathsf{M}_{K}^{\mathrm{spl}}$ and set $\mathcal{V}^{\mathrm{sub}}=\mathcal{V}^{\mathrm{can}}(-D^{\prime})$, which we call the sub-canonical extension of $\mathcal{V}=\mathcal{V}_{\eta}$.
@@
    $\widehat{\bigoplus}_{\ell\in\sigma^{\vee}}\Big{(}(\Psi^{\mathrm{spl}}_{Z}(\ell))^{\wedge}_{\overline{x}}\otimes\mathcal{V}_{0,\overline{x}}\Big{)},$
@@
One needs to check that in the smooth reduction case the sheaf $\mathcal{V}^{\mathrm{can}}$ satisfies the above two conditions. This can be achieved by modifying the arguments in the proof of [34] Proposition 5.6, where we take account of the splitting structures everywhere. ∎
@@ 5.4. Extensions of $\mathcal{G}_{0}^{\mathrm{spl}}$-zips and Hasse invariants
Before discussing extensions of zips to the boundary, we need some group theoretical preparation. For each $Z^{\mathrm{spl}}\subset\mathsf{M}^{\mathrm{spl},\mathrm{min}}_{K}$ and the associated $Z\subset\mathsf{M}^{\mathrm{min}}_{K}$, recall that we have the boundary PEL-type $\operatorname{\mathcal{O}}_{B}$-lattice $(\Lambda^{Z},\langle\cdot,\cdot\rangle^{Z},h^{Z})$. Over $\mathbb{Q}$, we have the associated parabolic subgroups
$Q=Q_{Z}\subset P=P_{Z}\subset G,$
cf. [38] Lemma 3.3.6 and Definition 3.3.8. More precisely, attached to the boundary stratum $Z\subset\mathsf{M}^{\mathrm{min}}_{K}$ we have a symplectic filtration $0\subset V_{-2}\subset V_{-1}\subset V_{0}=V$ satisfying the condition as in Lemma 3.3.6 of [38]. The group $P=P_{Z}$ is defined as the stabilizer of this filtration. The group $Q=Q_{Z}$ is defined as the kernel of the homomorphism $P\rightarrow\operatorname{GL}(gr_{-2}V)\times\operatorname{GL}(gr_{0}V),\;(g,r)\mapsto(r^{-1}gr_{-2}(g),gr_{0}(g))$, where $P\rightarrow\mathbb{G}_{m},\;(g,r)\mapsto r$ is the similitude character. The groups $P$ and $Q$ have the same unipotent radical $U$, such that $M_{h}:=Q/U$ is the reductive group for the rational PEL datum corresponding to $Z$, and $M:=P/U=M_{h}\times M_{l}$ where $M_{l}\simeq P/Q$ is a reductive group factor of the Levi $M$, see [38] Definition 3.3.8.
Consider the reductive group $\mathcal{G}_{0}^{\mathrm{spl}}$ over $\operatorname{\mathbb{F}}_{p}$. We want to adapt the construction of [2] subsection 3.3 to our setting. Consider the induced symplectic filtration
@@
For (1), we just repeat the construction in subsection 3.2 starting from $(\underline{\mathcal{H}}^{\mathrm{ext}},\underline{\operatorname{\mathscr{F}}}^{\mathrm{ext}},\underline{\iota}^{\mathrm{ext}})$ together with its universal splitting structure.
@@
is smooth. If $x\in\mathsf{M}^{\mathrm{spl}}_{K,0}$, this has been done in Proposition 3.10. So we may assume $x\in Z_{[\sigma]}^{\mathrm{spl}}$ for a boundary stratum. Then there is a corresponding point $x^{\prime}\in\Xi^{\mathrm{spl}}(\sigma)$ such that $C=\widehat{\operatorname{\mathcal{O}}}_{\Xi^{\mathrm{spl}}(\sigma),x^{\prime}}$. Let $y\in Z^{\mathrm{spl}}$ be the image of $x$ and $D=\widehat{\operatorname{\mathcal{O}}}_{Z^{\mathrm{spl}},y}$. We get an induced morphism
@@
To show the smoothness of $\zeta_{C}:\operatorname{Spec}\,C\rightarrow\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{\kappa}$, we proceed in four steps.
Step 1. We show $\zeta_{C}=\tau_{Z}\circ\zeta_{D}\circ\oint$. In other words, the following diagram commutes
${\operatorname{Spec}\,C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\oint}$$\scriptstyle{\zeta_{C}}$${\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}.}$${\operatorname{Spec}\,D\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\zeta_{D}}$${\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau_{Z}}$
@@
[…]
Step 2. $\oint:\operatorname{Spec}\,C\rightarrow\operatorname{Spec}\,D$ is smooth. For this, we apply the local diagrams for $\mathsf{M}^{\mathrm{spl},\mathrm{tor}}_{K,\Sigma}$ and $Z^{\mathrm{spl}}$ to realize $C$ and $D$ as local rings of the corresponding local models, which are also completions along identities of some unipotent subgroups, see the proof of Proposition 3.10. Then one can prove the smoothness of $\oint$ by checking that on tangent spaces it induces a projection.
Step 3. $\zeta_{D}:\operatorname{Spec}\,D\rightarrow\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}$ is smooth. As $Z^{\mathrm{spl}}$ is smooth, the smoothness of $\zeta_{D}$ follows from the boundary version of Proposition 3.10.
Step 4. $\tau_{Z}:\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}\rightarrow\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}$ is smooth. This follows from a similar argument as the proof of [2] Lemma 3.4. ∎
@@
The Hasse invariant $h_{w}\in H^{0}(\overline{X_{w}},\omega_{\operatorname{Hdg}}^{N_{w}})$ of Corollary 4.7 extends to a $G(\operatorname{\mathbb{A}}_{f}^{p})$-equivariant section $h_{w}^{\mathrm{tor}}\in H^{0}(\overline{X^{\mathrm{tor}}_{w}},\omega_{\operatorname{Hdg}}^{N_{w}})$ with non-vanishing locus being precisely $X_{w}^{\mathrm{tor}}$.
@@
Similar to [2] Corollary 3.7, this follows from the fact that for each boundary $Z$, $\tau_{Z}:\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}\rightarrow\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}$ is smooth, thus open, therefore its image does not contain the closed point of $\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}$. Indeed, if its image contained the closed point, it would be surjective and any maximal chain of points of $|\operatorname{\mathcal{G}}^{\mathrm{spl}}_{0}\textrm{-}\mathrm{Zip}^{\mu}_{k}|$ would be the image of a maximal chain of points of $|\mathcal{G}_{0,Z}^{\mathrm{spl}}\textrm{-Zip}^{\mu_{Z}}_{k}|$. This is impossible since $\dim\,Z<\dim\,X$ with $X=\mathsf{M}^{\mathrm{spl}}_{K,0}$ (see also the proof of [2] Corollary 3.7). ∎